a mother is three as old as her daughter six years from now,the mother will just be twice as old as her daughter.find their present ago
Answer by lwsshak3(11628)

a mother is three times as old as her daughter six years from now,the mother will just be twice as old as her daughter.find their present ages.
***
let x=daughter's age now
3x=mother's age now
..
6 yrs from now:
x+6=daughter's age
3x+6=mother's age
3x+6=2(x+6)
3x+6=2x+12
x=6
3x=18
daughter's present age=6
mother's present age=18